OpenStudy (anonymous):

Can someone explain how to multiply everything out? You should get 33j-198=27j from (j - 6)(11) = (j)(9). Rewarding medal and fan

OpenStudy (anonymous):

So we want to expand the brackets: 11(j-6) = (j)(9) 11*j - 11*6 = 9*j 11j - 66 = 9j Then you can times by 3 to get the answer you have given.

OpenStudy (anonymous):

why do I multiply by 3 after I cross multiply?

OpenStudy (fibonaccichick666):

It's not cross multiplying, it's called distributing. here is a reference http://www.purplemath.com/modules/simparen.htm

OpenStudy (fibonaccichick666):

and, just so you end up with the answer they want